# Evaluating Compact LLMs for Zero-Shot Iberian Language Tasks on End-User Devices

### v1

@@
Luís Couto Seller1, Íñigo Sanz Torres1, Adrián Vogel-Fernández1,
@@ 1 Introduction
The remainder of this article is structured as follows. Section 2 reviews related work on multilingual LLM evaluation. Section 3 introduces the LLMs selected for evaluation. Section 4 details the Iberian NLP tasks and datasets employed. Section 5 presents the experimental results, while Section 6 provides an analysis and discussion of these results. Finally, Section 7 summarizes the main findings and outlines future research directions.
@@ 2 Related Work
DS-R1-Qwen-7B - 0.505 - 0.542 0.616
Llama-3.1-8B 0.478 0.797 0.726 0.667 0.605
Ministral-3B 0.421 0.642 0.342 0.609 0.727
Ministral-8B 0.483 0.719 0.348 0.650 0.753
Mistral-7B 0.425 0.642 0.268 0.625 0.742
Qwen2.5-7B - 0.637 0.579 0.742 0.759
Gemma 2-9B 0.528 0.684 0.402 0.752 0.806
Salamandra-2B - 0.354 - - -
@@ 5 Results
This section presents the experimental outcomes of evaluating several compact LLMs on a diverse set of Iberian language tasks.
@@ 5.2.1 Reading Comprehension
DS-R1-Qwen-7B 0.516 0.610 0.293 0.538 0.647
Llama-3.1-8B 0.638 0.690 0.432 0.646 0.720
Ministral-3B 0.264 0.274 0.287 0.266 0.268
Ministral-8B 0.756 0.804 0.550 0.770 0.803
Mistral-7B 0.667 0.637 0.330 0.592 0.673
Qwen2.5-7B 0.799 0.833 0.462 0.800 0.853
Gemma 2-9B 0.894 0.892 0.790 0.878 0.903
Salamandra-2B 0.298 0.241 0.287 0.227 0.261
@@ 5.2.2 Translation and Adaptation
DS-R1-Qwen-7B 0.088 0.099 0.009 0.079 0.133
Llama-3.1-8B 0.260 0.211 0.090 0.232 0.277
Ministral-3B 0.001 0.001 $\approx$ 0 $\approx$ 0 0.001
Ministral-8B 0.231 0.185 0.073 0.220 0.273
Mistral-7B 0.241 0.190 0.024 0.172 0.248
Qwen2.5-7B 0.195 0.193 0.016 0.190 0.253
Gemma 2-9B 0.290 0.233 0.135 0.262 0.314
Salamandra-2B 0.159 0.130 0.072 0.146 0.156
@@
DS-R1-Qwen-7B 0.358 0.373 0.188 0.345 0.412
Llama-3.1-8B 0.543 0.510 0.385 0.521 0.557
Ministral-3B 0.071 0.080 0.061 0.068 0.074
Ministral-8B 0.506 0.470 0.351 0.506 0.546
Mistral-7B 0.518 0.484 0.233 0.467 0.524
Qwen2.5-7B 0.471 0.473 0.150 0.470 0.516
Gemma 2-9B 0.573 0.532 0.445 0.553 0.589
Salamandra-2B 0.454 0.424 0.393 0.448 0.445
@@
DS-R1-Qwen-7B 108.54 101.77 202.56 116.69 96.93
Llama-3.1-8B 67.43 71.08 112.66 71.43 64.36
Ministral-3B 485.44 448.44 516.05 507.63 483.41
Ministral-8B 72.24 75.26 114.62 72.71 64.39
Mistral-7B 71.30 75.73 143.43 79.53 69.10
Qwen2.5-7B 204.16 200.05 1181.98 201.60 189.45
Gemma 2-9B 60.05 65.98 84.57 62.59 57.46
Salamandra-2B 80.29 82.16 95.61 79.37 79.67
@@ 5.2.3 Math
DS-R1-Qwen-7B 0.032 0.110 0.040 0.012
Llama-3.1-8B 0.004 0.128 0.034 0.000
Ministral-3B 0.008 0.010 0.006 0.016
Ministral-8B 0.012 0.290 0.048 0.008
Mistral-7B 0.000 0.136 0.012 0.000
Qwen2.5-7B 0.096 0.390 0.042 0.100
Gemma 2-9B 0.032 0.368 0.000 0.000
Salamandra-2B 0.004 0.018 0.006 0.000
@@ 5.2.4 Question Answering
DS-R1-Qwen-7B 0.204 0.238 0.246
Llama-3.1-8B 0.282 0.352 0.294
Ministral-3B 0.216 0.198 0.202
Ministral-8B 0.280 0.360 0.308
Mistral-7B 0.284 0.330 0.272
Qwen2.5-7B 0.272 0.334 0.272
Gemma 2-9B 0.350 0.356 0.308
Salamandra-2B 0.290 0.310 0.252
@@ 5.2.5 Natural Language Inference
DS-R1-Qwen-7B 0.535 0.563 0.493
Llama-3.1-8B 0.437 0.493 0.563
Ministral-3B 0.535 0.479 0.563
Ministral-8B 0.437 0.634 0.437
Mistral-7B 0.507 0.592 0.592
Qwen2.5-7B 0.563 0.704 0.521
Gemma 2-9B 0.775 0.789 0.465
Salamandra-2B 0.563 0.451 0.437
@@
DS-R1-Qwen-7B 0.403 0.442 0.334
Llama-3.1-8B 0.513 0.494 0.411
Ministral-3B 0.335 0.336 0.336
Ministral-8B 0.527 0.466 0.425
Mistral-7B 0.508 0.439 0.315
Qwen2.5-7B 0.498 0.482 0.400
Gemma 2-9B 0.513 0.488 0.449
Salamandra-2B 0.486 0.466 0.430
@@ 5.2.6 Paraphrasing
DS-R1-Qwen-7B 0.599 0.550 0.551
Llama-3.1-8B 0.655 0.652 0.645
Ministral-3B 0.571 0.583 0.584
Ministral-8B 0.658 0.643 0.661
Mistral-7B 0.685 0.689 0.676
Qwen2.5-7B 0.664 0.651 0.605
Gemma 2-9B 0.679 0.669 0.629
Salamandra-2B 0.576 0.560 0.535
@@ 5.2.7 Irony and Sarcasm Detection
The irony and sarcasm detection task revealed clear differences in model performance across languages and architectures, as shown in Table 8. Some models consistently predicted the majority class (“No”), leading to misleadingly high accuracy due to the class imbalance. As a result, F1-score is reported instead, which better reflects performance across both classes. The best overall performance was achieved by Mistral Small 3-24B, which obtained the highest or near-highest F1 scores in all five languages, including 0.52 in Catalan and Spanish, and 0.53 in Galician. For Basque and Portuguese, its F1 scores remained high (0.48 and 0.51, respectively), indicating robust cross-lingual generalization in irony detection. Qwen2.5-7B and Gemma 2-9B also delivered strong results, with the latter showing highly consistent F1 scores (close to 0.49) in the five languages. Basque emerged as the most challenging language, with lower scores in most models, while Galician and Portuguese generally showed higher performance. These results emphasize the difficulty of irony detection and the value of larger, better-tuned models for achieving robust multilingual performance.
@@
DS-R1-Qwen-7B 0.616 0.463 0.121 0.362 0.182
Llama-3.1-8B 0.471 0.464 0.258 0.346 0.440
Ministral-3B 0.424 0.443 0.059 0.499 0.454
Ministral-8B 0.420 0.658 0.340 0.657 0.547
Mistral-7B 0.411 0.598 0.076 0.339 0.443
Qwen2.5-7B 0.411 0.544 0.621 0.546 0.589
Gemma 2-9B 0.45 0.356 0.544 0.372 0.510
Salamandra-2B 0.579 0.526 0.481 0.468 0.485
@@
DS-R1-Qwen-7B 0.150 0.338 0.031 0.457 0.244
Llama-3.1-8B 0.383 0.410 0.204 0.439 0.340
Ministral-3B 0.007 0.000 0.002 0.000 0.000
Ministral-8B 0.522 0.290 0.499 0.023 0.541
Mistral-7B 0.516 0.487 0.009 0.501 0.520
Qwen2.5-7B 0.490 0.500 0.117 0.430 0.460
Gemma 2-9B 0.467 0.466 0.490 0.485 0.493
Salamandra-2B 0.233 0.303 0.162 0.403 0.360
@@ 6.3 Execution Time Analysis
Table 9 reports the execution time for the reading comprehension tests (Belebele datasets). Smaller models, such as Mistral-3B and Salamandra-2B, completed the tests significantly faster. In contrast, larger models required more time to complete their tests, particularly Mistral Small 3-24B, which, even after quantization, took up to three times longer than the other models such as Gemma 2-9B or Llama3.1-8B.
Language-wise, the Basque tests took $\approx$7-15% longer than those in other languages for most models, except when using the Salamandra models, which were specifically trained on Iberian Languages.
@@
DS-R1-Qwen-7B 101 96 109 104 96
Llama-3.1-8B 102 96 111 99 95
Ministral-3B 56 56 65 57 56
Ministral-8B 94 92 100 94 93
Mistral-7B 102 104 117 105 102
Qwen2.5-7B 101 96 109 99 95
Gemma 2-9B 128 115 141 120 115
Salamandra-2B 42 42 41 42 42
@@ 6.4 Basque Analysis
The Basque language poses a unique challenge for NLP, as it is a language isolate with no known relatives. LLMs often leverage similarities among related languages, such as the Romance family, to generalize knowledge. In contrast, LLMs must learn Basque almost entirely from scratch, making the training process more demanding. ([Etxaniz et al., 2024] proposed a dedicated Basque corpus for training and evaluating language models, which was used to develop the model Latxa, a Llama-3.1-8B adaptation for Basque ([Etxaniz and others, 2024]. Figure 2 shows the performance comparison for Basque between Latxa and the strongest competing model in each task. Latxa underperforms in Reading Comprehension, Translation (evaluated using BLEU and ChrF), NLI on the WNLI benchmark and Irony Detection. However, it performs slightly better in Translation when evaluated with TER, as well as Math and NLI on the XNLI benchmark. Despite trained specifically for Basque, Latxa underperforms compared to similar size models, such as Ministral-8B. This may be partially attributed to the performance of its base model, which did not achieve top results across benchmarks, either in average multilingual performance or for Basque only. Still, Latxa improves upon Llama-3.1-8B, proving that specialized models iterated on monolingual data can achieve better performance.
@@ 7 Conclusions and Future Work
Future research should extend evaluations to new LLMs and additional tasks capturing subtler linguistic phenomena, such as sentiment and pragmatic reasoning. Efforts should also focus on improving performance for under‑resourced languages through targeted pretraining and fine‑tuning, and on refining deployment strategies for resource‑constrained environments.

### v2

@@
Luis Couto Seller1, Íñigo Sanz Torres1, Adrián Vogel-Fernández1,
@@ 2 Related Work
DS-R1-Qwen-7B-it - 0.505 - 0.542 0.616
Llama-3.1-8B-it 0.478 0.797 0.726 0.667 0.605
Ministral-3B-it 0.421 0.642 0.342 0.609 0.727
Ministral-8B-it 0.483 0.719 0.348 0.650 0.753
Mistral-7B-it 0.425 0.642 0.268 0.625 0.742
Qwen2.5-7B-it - 0.637 0.579 0.742 0.759
Gemma 2-9B-it 0.528 0.684 0.402 0.752 0.806
Salamandra-2B-it - 0.354 - - -
@@ 4.2 Irony/Sarcasm Benchmark
Literal translation may introduce model specific biases, but the homogeneous Iberian context limits divergences in irony expression, so we focus on performance comparison. A comprehensive bias analysis would require replicating sociodemographic polarization methods beyond the scope of this work ([Casola et al., 2024].
@@ 5 Results
This section presents the experimental outcomes of evaluating several compact LLMs on a diverse set of Iberian language tasks. Relevant scripts for data manipulation and evaluation are available at ([Funditec, 2025].
@@ 5.2.1 Reading Comprehension
DS-R1-Qwen-7B-it 0.516 0.610 0.293 0.538 0.647
Llama-3.1-8B-it 0.638 0.690 0.432 0.646 0.720
Ministral-3B-it 0.264 0.274 0.287 0.266 0.268
Ministral-8B-it 0.756 0.804 0.550 0.770 0.803
Mistral-7B-it 0.667 0.637 0.330 0.592 0.673
Qwen2.5-7B-it 0.799 0.833 0.462 0.800 0.853
Gemma 2-9B-it 0.894 0.892 0.790 0.878 0.903
Salamandra-2B-it 0.298 0.241 0.287 0.227 0.261
@@ 5.2.2 Translation and Adaptation
DS-R1-Qwen-7B-it 0.088 0.099 0.009 0.079 0.133
Llama-3.1-8B-it 0.260 0.211 0.090 0.232 0.277
Ministral-3B-it 0.001 0.001 $\approx$ 0 $\approx$ 0 0.001
Ministral-8B-it 0.231 0.185 0.073 0.220 0.273
Mistral-7B-it 0.241 0.190 0.024 0.172 0.248
Qwen2.5-7B-it 0.195 0.193 0.016 0.190 0.253
Gemma 2-9B-it 0.290 0.233 0.135 0.262 0.314
Salamandra-2B-it 0.159 0.130 0.072 0.146 0.156
@@
DS-R1-Qwen-7B-it 0.358 0.373 0.188 0.345 0.412
Llama-3.1-8B-it 0.543 0.510 0.385 0.521 0.557
Ministral-3B-it 0.071 0.080 0.061 0.068 0.074
Ministral-8B-it 0.506 0.470 0.351 0.506 0.546
Mistral-7B-it 0.518 0.484 0.233 0.467 0.524
Qwen2.5-7B-it 0.471 0.473 0.150 0.470 0.516
Gemma 2-9B-it 0.573 0.532 0.445 0.553 0.589
Salamandra-2B-it 0.454 0.424 0.393 0.448 0.445
@@
DS-R1-Qwen-7B-it 108.54 101.77 202.56 116.69 96.93
Llama-3.1-8B-it 67.43 71.08 112.66 71.43 64.36
Ministral-3B-it 485.44 448.44 516.05 507.63 483.41
Ministral-8B-it 72.24 75.26 114.62 72.71 64.39
Mistral-7B-it 71.30 75.73 143.43 79.53 69.10
Qwen2.5-7B-it 204.16 200.05 1181.98 201.60 189.45
Gemma 2-9B-it 60.05 65.98 84.57 62.59 57.46
Salamandra-2B-it 80.29 82.16 95.61 79.37 79.67
@@ 5.2.3 Math
DS-R1-Qwen-7B-it 0.032 0.110 0.040 0.012
Llama-3.1-8B-it 0.004 0.128 0.034 0.000
Ministral-3B-it 0.008 0.010 0.006 0.016
Ministral-8B-it 0.012 0.290 0.048 0.008
Mistral-7B-it 0.000 0.136 0.012 0.000
Qwen2.5-7B-it 0.096 0.390 0.042 0.100
Gemma 2-9B-it 0.032 0.368 0.000 0.000
Salamandra-2B-it 0.004 0.018 0.006 0.000
@@ 5.2.4 Question Answering
DS-R1-Qwen-7B-it 0.204 0.238 0.246
Llama-3.1-8B-it 0.282 0.352 0.294
Ministral-3B-it 0.216 0.198 0.202
Ministral-8B-it 0.280 0.360 0.308
Mistral-7B-it 0.284 0.330 0.272
Qwen2.5-7B-it 0.272 0.334 0.272
Gemma 2-9B-it 0.350 0.356 0.308
Salamandra-2B-it 0.290 0.310 0.252
@@ 5.2.5 Natural Language Inference
DS-R1-Qwen-7B-it 0.535 0.563 0.493
Llama-3.1-8B-it 0.437 0.493 0.563
Ministral-3B-it 0.535 0.479 0.563
Ministral-8B-it 0.437 0.634 0.437
Mistral-7B-it 0.507 0.592 0.592
Qwen2.5-7B-it 0.563 0.704 0.521
Gemma 2-9B-it 0.775 0.789 0.465
Salamandra-2B-it 0.563 0.451 0.437
@@
DS-R1-Qwen-7B-it 0.403 0.442 0.334
Llama-3.1-8B-it 0.513 0.494 0.411
Ministral-3B-it 0.335 0.336 0.336
Ministral-8B-it 0.527 0.466 0.425
Mistral-7B-it 0.508 0.439 0.315
Qwen2.5-7B-it 0.498 0.482 0.400
Gemma 2-9B-it 0.513 0.488 0.449
Salamandra-2B-it 0.486 0.466 0.430
@@ 5.2.6 Paraphrasing
DS-R1-Qwen-7B-it 0.599 0.550 0.551
Llama-3.1-8B-it 0.655 0.652 0.645
Ministral-3B-it 0.571 0.583 0.584
Ministral-8B-it 0.658 0.643 0.661
Mistral-7B-it 0.685 0.689 0.676
Qwen2.5-7B-it 0.664 0.651 0.605
Gemma 2-9B-it 0.679 0.669 0.629
Salamandra-2B-it 0.576 0.560 0.535
@@ 5.2.7 Irony and Sarcasm Detection
The irony and sarcasm detection task revealed clear differences in model performance across languages and architectures, as shown in Table 8. Some models consistently predicted the majority class (“No”), leading to misleadingly high accuracy due to the class imbalance. As a result, F1-score is reported instead, which better reflects performance across both classes. The best overall performance was achieved by Mistral Small 3-24B, which obtained the highest or near-highest F1 scores in all five languages, including 0.52 in Catalan and Spanish, and 0.53 in Galician. For Basque and Portuguese, its F1 scores remained high (0.48 and 0.51, respectively), indicating robust cross-lingual generalization in irony detection. Qwen2.5-7B and Gemma 2-9B also delivered strong results, with the latter showing highly consistent F1 scores (close to 0.49) in the five languages. Basque emerged as the most challenging language, with lower scores in most models, while Galician and Portuguese generally showed higher performance. Given that irony and sarcasm are deeply influenced by language and culture, the translated benchmark may not fully convey the intended meaning of each message. However, the consistent results across languages suggest the translations effectively preserved these nuances. These results emphasize the difficulty of irony detection and the value of larger, better-tuned models for achieving robust multilingual performance.
@@
DS-R1-Qwen-7B-it 0.616 0.463 0.121 0.362 0.182
Llama-3.1-8B-it 0.471 0.464 0.258 0.346 0.440
Ministral-3B-it 0.424 0.443 0.059 0.499 0.454
Ministral-8B-it 0.420 0.658 0.340 0.657 0.547
Mistral-7B-it 0.411 0.598 0.076 0.339 0.443
Qwen2.5-7B-it 0.411 0.544 0.621 0.546 0.589
Gemma 2-9B-it 0.45 0.356 0.544 0.372 0.510
Salamandra-2B-it 0.579 0.526 0.481 0.468 0.485
@@
DS-R1-Qwen-7B-it 0.150 0.338 0.031 0.457 0.244
Llama-3.1-8B-it 0.383 0.410 0.204 0.439 0.340
Ministral-3B-it 0.007 0.000 0.002 0.000 0.000
Ministral-8B-it 0.522 0.290 0.499 0.023 0.541
Mistral-7B-it 0.516 0.487 0.009 0.501 0.520
Qwen2.5-7B-it 0.490 0.500 0.117 0.430 0.460
Gemma 2-9B-it 0.467 0.466 0.490 0.485 0.493
Salamandra-2B-it 0.233 0.303 0.162 0.403 0.360
@@ 6.3 Execution Time Analysis
Table 9 reports the execution time for the reading comprehension tests (Belebele datasets). Smaller models, such as Mistral-3B and Salamandra-2B, completed the tests significantly faster. In contrast, larger models required more time to complete their tests. Particularly, Mistral Small 3-24B, even after quantization, took up to three times longer than other models such as Gemma 2-9B or Llama3.1-8B. Language-wise, Basque tests took $\approx$7-15% longer than those in other languages for most models, except when using Salamandra models, which were specifically trained on Iberian Languages.
@@
DS-R1-Qwen-7B-it 101 96 109 104 96
Llama-3.1-8B-it 102 96 111 99 95
Ministral-3B-it 56 56 65 57 56
Ministral-8B-it 94 92 100 94 93
Mistral-7B-it 102 104 117 105 102
Qwen2.5-7B-it 101 96 109 99 95
Gemma 2-9B-it 128 115 141 120 115
Salamandra-2B-it 42 42 41 42 42
@@ 6.4 Basque Analysis
The Basque language poses a unique challenge for NLP, as it is a language isolate with no known relatives. LLMs often leverage similarities among related languages, such as the Romance family, to generalize knowledge. In contrast, LLMs must learn Basque almost entirely from scratch, making the training process more demanding. ([Etxaniz et al., 2024] proposed a dedicated Basque corpus for training and evaluating language models, which was used to develop the model Latxa-Llama-3.1-8B-Instruct, a Llama-3.1-8B adaptation for Basque. Figure 2 shows the comparison for Basque between Latxa and the strongest competing model in each task. Latxa underperforms in Reading Comprehension, Translation (evaluated using BLEU and ChrF), NLI on the WNLI benchmark and Irony Detection. However, it performs slightly better in Translation when evaluated with TER, as well as Math and NLI on the XNLI benchmark.
Despite trained specifically for Basque, Latxa does not overperforms compared to similar size models, such as Gemma 2-9B or Ministral-8B. If all benchmark scores are normalized and averaged, Latxa is placed third after Mistral-3-24B and Gemma-2-9B. This may be partially attributed to the performance of its base model, which did not achieve top results across benchmarks, either in average multilingual performance or for Basque only. Still, Latxa-3.1-8B improves upon Llama-3.1-8B, proving that specialized models iterated on monolingual data can achieve better performance.
@@ 7 Conclusions and Future Work
Future research should extend evaluations to new LLMs and additional tasks capturing subtler linguistic phenomena, such as sentiment and pragmatic reasoning. Efforts should also focus on improving performance for under‑resourced languages through targeted pretraining and fine‑tuning, and on refining deployment strategies for resource‑constrained environments. Future work will also incorporate more statistically robust evaluation methods to ensure the reliability and significance of comparisons.